OpenStudy (anonymous):

solve on the interval 0 2pi 4 sec x+6=-2

OpenStudy (anonymous):

A. pi/6, 5pi/6 B. 2pi/3, 4pi/3 C. 7pi/6, 11pi/6 D. pi/3, 5pi/3 Help me please <3 :/

OpenStudy (anonymous):

\[4 \sec x+6=-2 \]\[\implies \sec x = -2 = \sec \frac \pi 3 \implies x = 2n \pi \pm \frac \pi 3, n\in \mathbb{W} \] Can you do from here?

OpenStudy (anonymous):

D.?

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

it was B

OpenStudy (anonymous):

Oops \[\sec x=-2=-\sec ( \frac \pi 3) = \sec (\pi -\frac \pi 3)= \sec \frac {2\pi} 3 \] now that becomes \[ \implies x = 2n \pi \pm \frac {2\pi} 3, n\in \mathbb{W} \] Hence B.